Thanks for all your help! :)
The heigh h, in meters, that Andreina is above the water t seconds after beginning her dive is given by the formula h = -5t^2 + 6t + 10. After how many seconds will Andreina hit the water, rounded to the nearest tenth of a second?

Answer by rothauserc(4718)

h = -5t^2 + 6t + 10
the problem asks what t(time) will be when h is 0
-5t^2 + 6t + 10 = 0
multiply both sides of the = by -
5t^2 -6t -10 = 0
use quadratic formula to solve
t = (6 + square root(36 + (4*5*10)) / 10) = 2.13622915
t = (6 - square root(36 + (4*5*10)) / 10) = −0.93622915
we want the positive value for t
t = 2.1 seconds